OpenStudy (anonymous):

A fire hose held near the ground shoots water at a speed of 6.5M/S . At what angle should the nozzle point in order that the water land 2.0M away?

OpenStudy (anonymous):

is this related with porjectile

OpenStudy (anonymous):

using special formula because it,s symetrical R=v^2sin 2(theta)/g so from here you can find (theta)

OpenStudy (anonymous):

\[\theta = \frac{ 1 }{2}\left( \sin^{-1} \frac{ 2*9.8 }{ \left( 6.5 \right)2 } \right)\]

OpenStudy (anonymous):

\[\theta = \sin^{-1} (x*g/v^{2})/2 \] theta is the angle of the nozzle x is horizontal displacement (2 meters) g is the gravity (9.8 meter/sec*sec) v is the initial velocity (6.5 meter/sec) theta=arcsin(2*9.8/6.5**2)/2 theta=0.24119926291439242 radians theta*180/pi = 13.819 degrees ! almost horizontal !